OpenStudy (anonymous):

what is 1/4 plus 11/12

OpenStudy (anonymous):

LOOK IT UP ON MATHWAY

OpenStudy (anonymous):

3/12+11/12=14/12=7/6

OpenStudy (anonymous):

i dont have that answer its answer choices

OpenStudy (unanimoose):

What are you choices? 7/6 should be correct.

OpenStudy (unanimoose):

You might have inserted the question wrong.

OpenStudy (anonymous):

i got 10/12,8/12,10/8,and 3/4

OpenStudy (anonymous):

\[\frac{ 1 }{ 4} + \frac{ 11 }{ 12 } =\frac{ 7 }{ 6 }\]

OpenStudy (unanimoose):

None of those are right, did you paste the question right?

OpenStudy (anonymous):

my question is 11/12-x=1/4

OpenStudy (anonymous):

So, are you solving for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[x = \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

i am 6th grade and i dont have that answer ugh

OpenStudy (anonymous):

I don't understand the question then. If you're solving for x then \[x = \frac{ 2 }{ 3 }\] would be the answer.

OpenStudy (anonymous):

well...

OpenStudy (anonymous):

sorry and thanks for trying

OpenStudy (anonymous):

My Pleasure. I'm In The 10th Grade. I'm Pretty Sure The Answer Would Be \[\frac{ 2 }{ 3 }\] If That Isn't The Answer Then Good Luck, I Have No Clue.

OpenStudy (anonymous):

\[\frac{11}{12}-x=\frac{1}{4}, x=\frac{2}{3} \]